OpenStudy (luigi0210):

Integration:

OpenStudy (luigi0210):

\[\LARGE \int_{-1}^{0}~\frac{e^{1/x}}{x^3}~dx\]

ganeshie8 (ganeshie8):

sub u = 1/x

myininaya (myininaya):

try letting u=1/x so x=1/u

OpenStudy (dan815):

then do parts

OpenStudy (anonymous):

Guys guys... In the eyes of god, mathematics doesn't exist! Now, let's all hold hands and pray.

ganeshie8 (ganeshie8):

yeah u can't avoid parts i think

OpenStudy (dan815):

mathematics is a language and God gave us languages

OpenStudy (dan815):

so shut your 3.14 hole

OpenStudy (anonymous):

LOLOL ^

OpenStudy (dan815):

:)

OpenStudy (luigi0210):

\[\LARGE \int_{-1}^{0}~\frac{e^{1/x}}{x^3}~dx\] \[\LARGE =\int_{-1}^{0}~\frac{1}{x}~e^{1/x}~\frac{1}{x^2}~dx\] \(\Large u=1/x\) and \(\Large du=-\frac{1}{x^2}dx\) \[\LARGE \int_{-1}^{0}~ue^udu\] And then use parts like dan said?

OpenStudy (luigi0210):

*\[\LARGE -\int_{-1}^{0}~ue^udu\]

myininaya (myininaya):

you should have an improper integral

OpenStudy (luigi0210):

\[\LARGE \LARGE \lim_{t \rightarrow 0} \int_{-1}^{t}~ue^udu~\]?

myininaya (myininaya):

of x is between -1 and 0 and u=1/x